A call center company has a total of 120 agents and 48 are men. what is the probability that a randomly selected call center agent from the company is a (a) women? (b) men?
Answer by macston(5194)

If there are 48 men, there are 120-48=72 women
.
a) 72/120=3/5
b) 48/120=2/5